# Incremental Model Order Reduction of Smoothed-Particle Hydrodynamic Simulations

### v1

@@ 1 Introduction
POD is generally considered a model order reduction approach that approximates a given dataset optimally on a low-dimensional basis [2]. Aiming at data compression, the strategy usually involves a Singular Value Decomposition (SVD) of a snapshot data matrix to derive a reduced basis spanned by the POD modes. Due to the often very good approximation by a small number of consecutive initial POD modes, only a few are usually considered at the expense of a small loss of information. Linearity assumptions limit the use of POD strategies. Nevertheless, the solution manifold is often sufficiently well approximated by a low-rank subspace, which recommends the use of the POD for model order reduction of nonlinear systems [37, 27, 25, 54, 55]. This is especially true for unsteady systems where the data is generated progressively in time, and the snapshot matrix usually consists of many more discrete spatial points (defining the length of the matrix column) than temporal points (defining the number of matrix columns). SVD procedures are traditionally offline approaches and thus require storing the entire snapshot matrix before performing the SVD, which still demands prohibitive computer memory. To solve this issue, efficient algorithmic alternatives dedicated to the incremental computation of the SVD (iSVD) that accompanies the data growth over time have been developed [16, 15, 5, 6, 26, 24].
@@
This paper investigates iSVD as a data compression technique for particle simulation methods. The goal is to reduce the storage requirements for future time-dependent, PDE-constrained optimization as much as possible while minimizing the required computational overhead. Though incrementally constructed ROMs have recently been used to handle memory limitations for optimization studies [50, 31, 32, 26, 29, 30], the iSVD has never been utilized for dynamic meshes and meshless methods such as Smoothed Particle Hydrodynamics (SPH). The challenges and innovations are considerable and relate to applying an iSVD to particle data obtained from massively parallel SPH simulations. Three aspects are particularly important: In contrast to grid-based methods, (a) the spatial positions of the data vary over time. Moreover, the snapshot matrix is irregularly populated because (b) the number of particles inside the domain usually varies over time, and (c) all particles are only active for a partial period when simulating open boundary domains.
@@ 3 Computational Model
The Lagrangian SPH method displays advantages over classical Eulerian flow simulation methods for problems with substantial temporal changes of the wetted domain. An example is violent free surface flows, such as those encountered in hydraulic machinery simulations. In these cases, grid-based methods often require dynamic grid refinement and coarsening, which is not trivial in 3D parallel frameworks. On the contrary, SPH approaches are mesh-free and discretize the material, i.e., it does not require the use or adaption of the computational grid to a free surface. Nowadays, SPH strategies are successfully applied to a broad range of engineering investigations, including –for example– free-surface flows [9], multi-continua applications [49], fluid-structure interaction [56], geotechnical [23], and turbine applications [34, 42]. Recent advancements include addressing volume conservation in sloshing flows and free-surface conditions [39], improving consistency and transport-velocity formulations via reverse kernel gradient correction [57], implicit shifting methods [41], Quasi-Lagrangian formulations, and energy balance techniques with Riemann solvers have also been introduced [35]. Results of the FOM employed in this work were obtained from the in-house Andritz solver, ASPHODEL. This section provides a brief overview of the implemented methods. A more detailed introduction can be found in [33, 28, 43, 38].
@@ 3.1.1 Smoothed-Particle Hydrodynamics Approximation
$\frac{dx_{i}}{dt}=v^{0}_{i}$ (9)
$\frac{d\omega_{i}}{dt}=\omega_{i}\sum_{j\in\mathcal{D}_{i}}\omega_{j}\left[\mathbf{v}^{0}_{j}-\mathbf{v}^{0}_{i}\right]\cdot\nabla_{i}W_{ij}$ (10)
$\frac{d(\omega_{i}\rho_{i})}{dt}=-\omega_{i}\sum_{j\in\mathcal{D}_{i}}2\omega_{j}\rho_{ij}^{E}\left[\mathbf{v}_{ij}^{E}-\mathbf{v}^{0}_{ij}\right]\cdot\nabla W_{ij}$ (11)
$\frac{d(\omega_{i}\rho_{i}\mathbf{v}_{i})}{dt}=\omega_{i}\rho_{i}\mathbf{g}-\omega_{i}\sum_{j\in\mathcal{D}_{i}}2\omega_{j}\left\{\rho_{ij}^{E}\mathbf{v}_{ij}^{E}\otimes[\mathbf{v}_{ij}^{E}-\mathbf{v}^{0}_{ij}]+p_{ij}^{E}\mathbf{I}\right\}\cdot\nabla_{i}W_{ij}\,.$ (12)
For simplicity, the system of equations (9-12) is presented without the inclusion of boundary terms in Eq. 8. Mind that the space conservation law (10) has been augmented by the subtraction of a term proportional to $\sum\nabla_{i}W_{ij}$ on the right-hand side (r.h.s.), which vanishes in continuous space, to improve the consistency of the discrete model. Similarly, the continuity equation (11) and the momentum equation (12) have been augmented by the addition of a term proportional to $\sum\nabla_{i}W_{ij}$ on the r.h.s., to support the conservation on the particle level. The superscript $E$ indicates interaction values between pairs of neighboring particles obtained by the solution of a Riemann problem, viz.
$\mathbf{v}_{ij}^{E}=\frac{\mathbf{v}_{i}+\mathbf{v}_{j}}{2}+\frac{p_{j}-p_{i}}{2\bar{\rho}_{ij}\bar{c}_{ij}},$ (6a)
$p_{ij}^{E}=\frac{p_{i}+p_{j}}{2}+\frac{\bar{\rho}_{ij}\bar{c}_{ij}(\mathbf{v}_{j}-\mathbf{v}_{i})}{2},$ (6b)
@@
The normalization factor $\sigma$ is given by $\sigma=\frac{3}{4\pi}$ in 2D and $\sigma=\frac{165}{256\pi}$ in 3D.
@@ 4.1 Singular Value Decomposition
The truncated SVD is the starting point of many algorithms utilized for dimensionality reduction, data compression, and noise reduction [48, 7]. Nevertheless, since the degree of achievable compression depends on the specific data, no a-priori criteria are usually available to decide how many modes are necessary to keep. Commonly used methods for truncation include setting a hard threshold for singular values, retaining a specified percentage of the total energy, or analyzing the singular values to identify an ”elbow” or ”knee” in their distribution [19, 18, 7]. It was suggested in [26] that ignoring the first $o$ singular values might be beneficial while computing the retained energy to avoid machine accuracy issues.
@@ 4.2 Incremental Singular Value Decomposition
Let us assume that $Y=USV^{T}$ is known, and that new data becomes available. The original matrix $Y$ and the updated matrix $\tilde{Y}$ are given by Eqn. 17 and Eqn. 18, respectively. The objective is to compute the SVD of $\tilde{Y}$, without having to reconstruct $Y$ first.
@@ 5 Model Reduction of SPH-Data
In the SPH framework, unique complexities arise from potential particle (a) deletion, (b) domain entering, or (c) exiting. In the latter case, the position of the calculation points remains unknown, resulting in additional fields –$\{x,y,z\}$– to reduce and store via SVD. Similarly, particles may lack data prior to their injection or after their deletion, leading to bunch matrices of irregular row and column size.
@@
We address missing data within $\mathbf{B}$ using one of the following approaches: imputation using the temporal mean of the available data (referred to here as Mean Imputation), imputation using the mean of subsets of the matrix defined by blocks of a given number of columns (referred to as Block-Mean Imputation), Gappy Proper Orthogonal Decomposition (GPOD), or a combination of Mean and Block-Mean Imputation. The GPOD is typically used to provide approximations for missing or corrupted data, cf. exemplary applications regarding noisy PIV measurement [44], inverse design [8], unsteady flow estimation [53], design optimization [11], or flows with deforming meshes [17]. However, since the missing values in SPH data refer to time instants when a particle is not in the considered domain, the approximated value has no physical meaning and only completes the matrix $B$ to compute the SVD. The only desired property of the imputed values is that they should not raise the rank of the matrix so that a sound reduction can be achieved. The basic idea behind the GPOD is to use the SVD to improve the guessed data for the missing values, initially applied by [13] and later modified by [21, 40]. According to Gunes algorithm for GPOD [21, 40, 44]:
@@
    The temporal mean of the available data is used as an initial guess to replace the gaps.
@@
    The gaps in the original data are filled using values from reconstructing the matrix with truncated SVD; the original data remains unchanged.
@@
A particle may leave the domain before the new bunch matrix is collected; in that case, the temporal mean is replaced by the spatial mean for all considered imputation strategies.
Particles injected past the last update will also miss left-singular vectors. One approach could be to divide the bunch-matrix into $B\in\mathbb{R}^{m_{1}\times b}$ to add the new columns and $C\in\mathbb{R}^{m_{2}\times(T+b)}$ to add the new rows, which is compatible with the algorithm in Sec. 4.2. Adding rows, which corresponds to adding $C$ to the SVD of $Y$, is analogous to adding the columns of $C^{\top}$ to $Y^{\top}=VSU^{\top}$. However, in addition to computing the SVD of $K_{B}\in\mathbb{R}^{(q+b)\times(q+b)}$, this requires a second SVD on $K_{C}\in\mathbb{R}^{(q+m_{2})\times(q+m_{2})}$. The number of particles injected between updates is often immense. Considering that SVD’s need to be computed for each $b$ time steps and their computational effort usually scales with $O(mn\min(m,n))$, the additional efforts of $K_{C}$ can become particularly costly. We decided to substitute the missing left-singular vectors for the $m_{2}$ new particles with the left singular vectors from other particles. This does not affect the singular values as well as the singular vectors of the existing SVD. One can use any row of $U$. For simplicity, when $m_{2}<m_{1}$ –which is the most common case–, the bottom $m_{2}$ rows of $U$ are utilized:
@@
If the data matrix $Y$ contains more than one field, the rows to be used depend on how the data has been organized when the snapshots were collected. However, for the remainder of this work, each field undergoes its own SVD and all fields are treated independently.
@@ 5.1 Parallelization
4:      $Q_{2},R=\text{serialQR}(\tilde{R})$ // Serial QR decomposition on gathered $R$
@@
The procedure to update the SVD is presented in Alg. 4 and follows [26] with some key differences: (a) Handling missing rows in $U$, (b) performing the initial SVD on a complete bunch-matrix rather than a single column vector, (c) neglecting the specification of a truncation criterion, and (d) using the TSQR decomposition instead of a modified Gram-Schmidt procedure. The algorithm starts by checking for a possibly empty local matrix $U$ – which would happen if enough particles were injected since the last update to activate a new process. In that case, the local $U$ is initialized as a zero matrix instead of copying left-singular vectors from other processes, which would require additional communication effort. The remaining cases with already populated local left-singular vectors extend $U$ depending on whether the number of additional rows exceeds the number of existing rows. The global product $U^{T}B$ is computed in serial mode and then made global by parallel summation, cf. [26]. The residual matrix $P=B-UM$ is orthogonalized using the proposed TSQR decomposition strategy from Alg 3. An additional orthonormalization step is applied to $Q$, which accounts for a potential numerical loss of orthogonality during SVD updates [6, 16, 14, 15, 1, 26]. Following the enhanced itSVD algorithm outlined in [26], this step precedes the construction of $K$. The global matrix $K$ is constructed and decomposed on one specific process to avoid rounding errors in the singular values. Truncated SVD results are then made global through broadcasting, and the SVD matrices are updated.
@@
18:$P=B-UM$
19:$Q_{p},R_{p}=\text{TSQR}(P)$
20:$Q=\begin{bmatrix}U&Q_{p}\end{bmatrix}$
21:$Q_{q},R_{q}=\text{TSQR}(Q)$
@@
32:$Bcast(U^{\prime},S,V^{\prime})$ // Broadcast truncated results to all processes
33:$R=\begin{bmatrix}V&0\\ 0&I\end{bmatrix}$
34:$U=Q_{q}U^{\prime}$
35:$V=RV^{\prime}$
@@ 6 Validation
[…]
@@
The 2D dam break simulation models the temporal evolution of an initially rectangular column of water ($\rho_{0}=1000$ kg/m3) under the influence of pressure and gravity. The complex sloshing motion involves different wave patterns and splashing as the water interacts with the walls. The test case is, therefore, frequently used to scrutinize SPH methods. The considered domain refers to a closed box, cf. Fig. 2, i.e., no particle can enter and leave the domain. This allows for an assessment of the incremental SVD, particularly its ability to capture the particle positions without entering missing data.
[…]
Figure 2: 2D Dam Break: Illustration of the initial configuration.
@@
In order to ensure efficiency, the adaptive truncation variant is preferred. When attention is directed to the reconstruction of the density and positions, cf. Figs. 4-6, it appears that the adaptive truncation essentially corresponds to a truncation at $q=50$, essentially confirming the results of Fig. 3. Except for very few outliers, maximum relative error magnitudes below one per cent are achievable with the adaptive scheme. The picture changes for the reconstruction of the velocities, displayed in Figs. 7-8, where the adaptive truncation agrees with the fixed-rank truncation at $q=200$ or $q=400$ for a similar error level, which reveals more complex dynamics of the velocities and a smaller degree of compression. The large discrepancy of the required rank recommends separate model reductions for each field considered to increase the efficiency.
Due to the non-periodic behavior of the dam break flow, more information is continuously added to the SVD as the simulation progresses, and the truncation ranks for each field keep growing, cf. Fig. 9a. Having multiple SVDs that handle separate time windows might help mitigate this problem. Although the segmented SVD approach, in combination with adaptive truncation, always shows a better agreement between FOM and ROM, there may be doubts about the associated data effort. For the example employed above, i.e., ten vs. one window, Fig. 9b reveals an average maximum rank for the vertical velocity $w$ above 100. When multiplied by ten windows, this is more than the maximum rank of the single-window setup indicated in Fig. 9a, which is slightly below 800 and would give the impression that the 10-window setup is less efficient than the single-window setup. However, the associated data effort for multiple windows is estimated by more than simply adding up the maximum rank for all windows. Considering that the size of the left singular vector is determined by the product of the rank and the number of particles, the size of the singular values corresponds to the rank. The product of the number of time steps and the rank determines the size of the right singular vector. The following equation therefore gives the compression ratio for the windowing strategy $CR_{w}$:
@@
[…]
@@
All four imputation strategies described in Sec. 5, i.e. mean (MI), block-mean (BMI), hybrid MI-BMI and gappy POD (GPOD) methods. The hybrid mean and block-mean imputation, denoted as $h_{MI-BMI}$, refers to a situation, where a bunch matrix with a standard deviation below unity after the MI is replaced by BMI. Furthermore, the GPOD approach is assessed for three different (fixed) iteration numbers, i.e., GPOD(i), where $i\in[1,5,10]$.
Figure 11 displays an exemplary comparison of the reconstructed jet and the full-order SPH data extracted at three different time instants, with $b=250$.

### v2

@@ 1 Introduction
POD is generally considered a model order reduction approach that approximates a given dataset optimally on a low-dimensional basis [5]. Aiming at data compression, the strategy usually involves a Singular Value Decomposition (SVD) of a snapshot data matrix to derive a reduced basis spanned by the POD modes. Due to the often very good approximation by a small number of consecutive initial POD modes, only a few are usually considered at the expense of a small loss of information. Linearity assumptions limit the use of POD strategies. Nevertheless, the solution manifold is often sufficiently well approximated by a low-rank subspace, which recommends the use of the POD for model order reduction of nonlinear systems [46, 35, 32, 65, 66, 1, 4]. This is especially true for unsteady systems where the data is generated progressively in time, and the snapshot matrix usually consists of many more discrete spatial points (defining the length of the matrix column) than temporal points (defining the number of matrix columns). SVD procedures are traditionally offline approaches and thus require storing the entire snapshot matrix before performing the SVD, which still demands prohibitive computer memory. To solve this issue, efficient algorithmic alternatives dedicated to the incremental computation of the SVD (iSVD) that accompanies the data growth over time have been developed [22, 21, 8, 9, 34, 33].
@@
This paper investigates iSVD as a data compression technique for particle simulation methods. The goal is to reduce the storage requirements for future time-dependent, PDE-constrained optimization as much as possible while minimizing the required computational overhead. Though incrementally constructed ROMs have recently been used to handle memory limitations for optimization studies [61, 40, 41, 34, 37, 38], the iSVD has never been utilized for dynamic meshes and meshless methods such as Smoothed Particle Hydrodynamics (SPH). The challenges and innovations are considerable and relate to applying an iSVD to particle data obtained from massively parallel SPH simulations. Three aspects are particularly important. In contrast to grid-based methods, (a) the spatial positions of the data vary over time; (b) the snapshot matrix is irregularly populated because the number of particles within the domain typically varies over time; (c) all particles are only active during a partial period when simulating open boundary domains.
@@ 3 Computational Model
The Lagrangian SPH method displays advantages over classical Eulerian flow simulation methods for problems with substantial temporal changes of the wetted domain. An example is violent free surface flows, such as those encountered in hydraulic machinery simulations. In these cases, grid-based methods often require dynamic grid refinement and coarsening, which is not trivial in 3D parallel frameworks. On the contrary, SPH approaches are mesh-free and discretize the material, i.e., it does not require the use or adaption of the computational grid to a free surface. Nowadays, SPH strategies are successfully applied to a broad range of engineering investigations, including –for example– free-surface flows [13, 71, 18], multi-continua applications [60], fluid-structure interaction [69, 14, 54], geotechnical [29, 39], energy harvesting from water waves and wind offshore [67, 12], and turbine applications [43, 51]. Recent advancements include addressing volume conservation in sloshing flows and free-surface conditions [48], improving consistency and transport-velocity formulations via reverse kernel gradient correction [70], implicit shifting methods [50], Quasi-Lagrangian formulations, and energy balance techniques with Riemann solvers have also been introduced [44]. Results of the FOM employed in this work were obtained from the in-house Andritz solver, ASPHODEL. This section provides a brief overview of the implemented methods. A more detailed introduction can be found in [42, 36, 52, 47].
@@ 3.1.1 Smoothed-Particle Hydrodynamics Approximation
$\frac{d\mathbf{x}_{i}}{dt}=\mathbf{v}^{0}_{i}$ (9)
$\frac{d\omega_{i}}{dt}=\omega_{i}\sum_{j\in\mathcal{D}_{i}}\omega_{j}\left[\mathbf{v}^{0}_{j}-\mathbf{v}^{0}_{i}\right]\cdot\boldsymbol{\nabla}_{i}W_{ij}$ (10)
$\frac{d(\omega_{i}\rho_{i})}{dt}=-\omega_{i}\sum_{j\in\mathcal{D}_{i}}2\omega_{j}\rho_{ij}^{E}\left[\mathbf{v}_{ij}^{E}-\mathbf{v}^{0}_{ij}\right]\cdot\boldsymbol{\nabla}_{i}W_{ij}$ (11)
$\frac{d(\omega_{i}\rho_{i}\mathbf{v}_{i})}{dt}=\omega_{i}\rho_{i}\mathbf{g}-\omega_{i}\sum_{j\in\mathcal{D}_{i}}2\omega_{j}\left\{\rho_{ij}^{E}\mathbf{v}_{ij}^{E}\otimes[\mathbf{v}_{ij}^{E}-\mathbf{v}^{0}_{ij}]+p_{ij}^{E}\mathbf{I}\right\}\cdot\boldsymbol{\nabla}_{i}W_{ij}\,.$ (12)
For simplicity, the system of equations (9-12) is presented without the inclusion of boundary terms in Eq. 8. Mind that the space conservation law (10) has been augmented by the subtraction of a term proportional to $\sum\boldsymbol{\nabla}_{i}W_{ij}$ on the right-hand side (r.h.s.), which vanishes in continuous space, to improve the consistency of the discrete model. Similarly, the continuity equation (11) and the momentum equation (12) have been augmented by the addition of a term proportional to $\sum\boldsymbol{\nabla}_{i}W_{ij}$ on the r.h.s., to support the conservation on the particle level. The superscript $E$ indicates interaction values between pairs of neighboring particles obtained by the solution of a Riemann problem, viz.
$\mathbf{v}_{ij}^{E}=\frac{\mathbf{v}_{i}+\mathbf{v}_{j}}{2}+\frac{p_{j}-p_{i}}{2\bar{\rho}_{ij}\bar{c}_{ij}},$ (13)
$p_{ij}^{E}=\frac{p_{i}+p_{j}}{2}+\frac{\bar{\rho}_{ij}\bar{c}_{ij}(\mathbf{v}_{j}-\mathbf{v}_{i})}{2},$ (14)
@@
Here, $W(\mathbf{x},h)$ represents the smoothing kernel function, $\mathbf{x}$ is the distance vector between two particles, and $h$ is the smoothing length. The parameter $d$ represents the number of spatial dimensions in the simulation, $\|\mathbf{x}\|$ denotes the Euclidean norm of the distance vector $\mathbf{x}$. The function $\theta(q)$ defines the shape of the kernel as a function of the normalized distance $q=\|\mathbf{x}\|/(2h)$. The subscript $()_{+}$ represents the positive part, meaning that the function goes to zero if the input is negative. The normalization factor $\sigma$ is given by $\sigma=\frac{3}{4\pi}$ in 2D and $\sigma=\frac{165}{256\pi}$ in 3D.
@@ 4.1 Singular Value Decomposition
The truncated SVD is the starting point of many algorithms utilized for dimensionality reduction, data compression, noise reduction, solving systems of algebraic equations, inverting rank-deficient matrices, and more [58, 10, 30, 31, 59]. Nevertheless, since the degree of achievable compression depends on the specific data, no a-priori criteria are usually available to decide how many modes are necessary to keep. Commonly used methods for truncation include setting a hard threshold for singular values, retaining a specified percentage of the total energy, or analyzing the singular values to identify an ”elbow” or ”knee” in their distribution [25, 24, 10]. It was suggested in [34] that ignoring the first $o$ singular values might be beneficial while computing the retained energy to avoid machine accuracy issues.
@@ 4.2 Incremental Singular Value Decomposition
Let us assume that $Y=USV^{T}$ is known, and that new data becomes available. The original matrix $Y$ and the updated matrix $\tilde{Y}$ are given by Eqn. 19 and Eqn. 20, respectively. The objective is to compute the SVD of $\tilde{Y}$, without having to reconstruct $Y$ first.
@@ 5 Model Reduction of SPH-Data
In the SPH framework, unique complexities arise from potential particle (a) deletion, (b) domain entering, or (c) exiting. In the latter case, the position of the calculation points remains unknown, resulting in additional fields $\{x,y,z\}$– to reduce and store via SVD. Similarly, particles may lack data prior to their injection or after their deletion, leading to bunch matrices of irregular row and column size.
@@
here, the entries $\mathbf{y}_{p,t}$ refer to time instants $t$ when a particle $p$ was present in the domain and are referred to herein as available data. We address missing data within $\mathbf{B}$ using one of the following approaches: if at least one non-null value exists for particle $p$ within $\mathbf{B}$, the initial guess for its value at time $t$ is computed as the row mean of all available non-null entries for particle $p$ in $\mathbf{B}$; if no non-null values are available for particle $p$ in the current $\mathbf{B}$, then for each column $t$, the initial guess is set to the mean of the non-null values in that column (referred to here as Mean Imputation), imputation using the mean of subsets of the matrix defined by blocks of a given number of columns (referred to as Block-Mean Imputation), Gappy Proper Orthogonal Decomposition (GPOD), or a combination of Mean and Block-Mean Imputation. The GPOD is typically used to provide approximations for missing or corrupted data, cf. exemplary applications regarding noisy PIV measurement [53], inverse design [11], unsteady flow estimation [64], design optimization [16], or flows with deforming meshes [23]. However, since the missing values in SPH data refer to time instants when a particle is not in the considered domain, the approximated value has no physical meaning and only completes the matrix $B$ to compute the SVD. The only desired property of the imputed values is that they should not raise the rank of the matrix so that a sound reduction can be achieved. The basic idea behind the GPOD is to use the SVD to improve the guessed data for the missing values, initially applied by [19] and later modified by [27, 49]. According to Gunes algorithm for GPOD [27, 49, 53]:
@@
    The temporal mean of the available data within $\mathbf{B}$ is used as an initial guess to replace the gaps.
@@
    The gaps within the original $\mathbf{B}$ are filled using values from reconstructing the truncated SVD of $\mathbf{B}$ ; the original data remains unchanged.
@@
Particles injected past the last update will also miss left-singular vectors. One approach could be to divide the bunch-matrix into $B\in\mathbb{R}^{m_{1}\times b}$ to add the new columns and $C\in\mathbb{R}^{m_{2}\times(T+b)}$ to add the new rows, which is compatible with the algorithm in Sec. 4.2. Adding rows, which corresponds to adding $C$ to the SVD of $Y$, is analogous to adding the columns of $C^{\top}$ to $Y^{\top}=VSU^{\top}$. However, in addition to computing the SVD of $K_{B}\in\mathbb{R}^{(q+b)\times(q+b)}$, this requires a second SVD on $K_{C}\in\mathbb{R}^{(q+m_{2})\times(q+m_{2})}$. The number of particles injected between updates is often immense. Considering that SVD’s need to be computed for each $b$ time steps and their computational effort usually scales with $O(mn\min(m,n))$, the additional efforts of $K_{C}$ can become particularly costly. We decided to substitute the missing $U$-rows for the $m_{2}$ new particles with existing $U$-rows. This does not affect the singular values as well as the singular vectors of the existing SVD. One can use any row of $U$. For simplicity, when $m_{2}<m_{1}$ –which is the most common case–, the bottom $m_{2}$ rows of $U$ are utilized:
@@
Mind that adding rows to $\mathbf{U}$ is analogous to adding rows to $\mathbf{Y}$. Opting to replace the missing rows with existing rows, the data prior to a particle’s injection are copies of data related to particles that were already active at that time instant of the simulation. Mind that these values are not physically meaningful and are discarded during post-processing since the new particle was not active yet. If the data matrix $Y$ contains more than one field, the rows to be used depend on how the data has been organized when the snapshots were collected. However, for the remainder of this work, each field undergoes its own SVD and all fields are treated independently.
@@ 5.1 Parallelization
5:    $\tilde{Q}=Q_{2}[\text{procId}\cdot A.\text{cols()}:(\text{procId}+1)\cdot A.\text{cols()},:]$ // Extract process specific data
@@
The procedure to update the SVD is presented in Alg. 4 and follows [34] with some key differences: (a) Handling missing rows in $U$, (b) performing the initial SVD on a complete bunch-matrix rather than a single column vector, (c) neglecting the specification of a truncation criterion, and (d) using the TSQR decomposition instead of a modified Gram-Schmidt procedure. The algorithm starts by checking for a possibly empty local matrix $U$ – which would happen if enough particles were injected since the last update to activate a new process. In that case, the local $U$ is initialized as a zero matrix instead of copying left-singular vectors from other processes, which would require additional communication effort. The remaining cases with already populated local left-singular vectors extend $U$ depending on whether the number of additional rows exceeds the number of existing rows. The global product $U^{T}B$ is computed in serial mode and then made global by parallel summation, cf. [34]. The residual matrix P=BUMP=B-UM is orthogonalized using the proposed TSQR decomposition strategy from Alg 3. An additional orthonormalization step is applied to $Q$, which accounts for a potential numerical loss of orthogonality during SVD updates [9, 22, 20, 21, 3, 34]. Following the enhanced itSVD algorithm outlined in [34], this step precedes the construction of $K$. The global matrix $K$ is constructed and decomposed on one specific process to avoid rounding errors in the singular values. Truncated SVD results are then made global through broadcasting, and the SVD matrices are updated.
@@
18:P=BUMP=B-UM
[…]
@@
[…]
@@ 6 Validation
which observes the maximum difference between Full Order Model (SPH) and Reduced Order Model (SVD) for a field $\phi$ at time $\tau$, with respect to the maximum absolute value during the complete simulation $\phi_{\text{ref}}$, where $\|\cdot\|$ represents the absolute value. The SVD is either adaptively truncated by neglecting all singular values that are five orders of magnitude smaller than the first singular value or always truncated at a fixed rank. Secondly, the overhead of the procedures is assessed, which involves both the computational time and the associated storage/memory efforts. For short-term dynamics with low periodicity and limited temporal correlations, a segmentation of the time window under consideration for the SVD could be advantageous. The latter is associated with separate ROMs for each temporal segment combined with adaptive truncation criteria. Related computational advantages are, therefore, not guaranteed and will be investigated in the first case of this study. In this context, we will also examine the changes in the reconstruction error when moving from one segment to another.
All simulations were conducted on an AMD EPYC 7573X 32-Core Processor (2 sockets), 512GiB of RAM, and an NVIDIA RTX A6000 GPU (48GiB memory) for the 3D application.
@@
The 2D dam break simulation models the temporal evolution of an initially rectangular column of water ($\rho_{0}=1000$ kg/m3) under the influence of pressure and gravity. The complex sloshing motion involves different wave patterns and splashing as the water interacts with the walls. The test case is, therefore, frequently used to scrutinize SPH methods, see [2, 68]. The considered domain refers to a closed box, cf. Fig. 2, i.e., no particle can enter and leave the domain. This allows for an assessment of the incremental SVD, particularly its ability to capture the particle positions without entering missing data.
The square domain spans $2H\times 2H$, where $H=2m$ denotes the initial water column height. The initial width of the water column refers to $H/2$, and the initial particle spacing is assigned to $\Delta x=0.02m$, which amounts to 5000 fluid particles. The simulation is performed over 20 000 time steps and terminates before the flow comes to rest, i.e., after the waterfront hits the downstream wall, goes up, and then comes back down. The Courant number criterion (18) controls the time step. To preserve a small Mach number $M\leq 0.1$, the numerical speed of sound is assigned to $c_{0}=100\,\text{m/s}$ which is well above an estimated Torricelli speed of $\sqrt{2\,g\,H}\approx 6.3m/s$, with $\mathbf{g}=-g\mathbf{e}_{z}$ and $g=9.81m/s^{2}$. Mind that the bunch size used for the SVD updates is assigned to $b=400$ in this case.
Figure 2: 2D Dam Break: Illustration of the initial configuration. Three height measurement probes (X1, X2, X3) are highlighted in red.
@@
In order to ensure efficiency, the adaptive truncation variant is preferred. When attention is directed to the reconstruction of the density and positions, cf. Figs. 4-6, it appears that the adaptive truncation essentially corresponds to a truncation at $q=50$, essentially confirming the results of Fig. 3. Except for very few outliers, maximum relative error magnitudes below one per cent are achievable with the adaptive scheme. The picture changes for the reconstruction of the velocities, displayed in Figs. 7-8, where the adaptive truncation agrees with the fixed-rank truncation at $q=200$ or $q=400$ for a similar error level, which reveals more complex dynamics of the velocities and a smaller degree of compression. The large discrepancy of the required rank recommends separate model reductions for each field considered to increase the efficiency. Fig. 9 depicts the temporal evolution of the computed water level (height) at the locations X1, X2, and X3, cf. Fig. 2. The figure shows excellent agreement between the FOM and the ROM results. Note that isolated splashes and droplets were not taken into account when measuring the height.
Due to the non-periodic behavior of the dam break flow, more information is continuously added to the SVD as the simulation progresses, and the truncation ranks for each field keep growing, cf. Fig. 10a. Having multiple SVDs that handle separate time windows might help mitigate this problem. Let us consider two exemplary setups: one option in which a single SVD covers all 20,000 time steps, and a second option in which 10 separate SVDs each cover 2,000 non-overlapping time steps. Both setups feature the same bunch width, i.e. $b=400$. Although the segmented SVD approach, in combination with adaptive truncation, always shows a better agreement between FOM and ROM, there may be doubts about the associated data effort. For the example employed above, i.e., ten vs. one window, Fig. 10b reveals an average maximum rank for the vertical velocity $w$ above 100. When multiplied by ten windows, this is more than the maximum rank of the single-window setup indicated in Fig. 10a, which is slightly below 800 and would give the impression that the 10-window setup is less efficient than the single-window setup. However, the associated data effort for multiple windows is estimated by more than simply adding up the maximum rank for all windows. Considering that the size of the left singular vector is determined by the product of the rank and the number of particles, the size of the singular values corresponds to the rank. The product of the number of time steps and the rank determines the size of the right singular vector. The following equation therefore gives the compression ratio for the windowing strategy $CR_{w}$:
@@
Figure 9: 2D Dam Break: Temporal evolution of the water level (height) at the horizontal locations X1, X2, and X3, obtained from the ROM with adaptive truncation. Left graph shows single window/SVD results and right graph displays results of a segmentation into 10 (equally sized) windows/SVDs, both in combination with a bunch matrix width $b=400$.
@@
The second test case concerns the simulation of an impinging jet. The rectangular domain has a horizontal bottom boundary corresponding to a flat plate with a length $L_{p}=0.18$m, and the height of the domain reads H=0.1H=0.1m. Particles are injected vertically through a confined, centered area of width L=0.02L=0.02m from the top with a velocity $w_{in}=-20\,\text{m/s}$ and flow towards the plate along its normal. Figure 11a illustrates the initial particle configuration that employs a regular Cartesian lattice distance $\Delta x=0.0005$ and amounts to 8000 particles. The speed of sound is assigned to $c_{0}=200\,\text{m/s}$, i.e. $w_{in}/c_{0}=0.1$, and the influence of gravity is neglected.
The simulation is performed over $N_{s}=$ 5 000 time steps and terminates in a fully developed state illustrated in Fig. 11b. The total number of active particles in the fully-developed (pseudo-steady) state oscillates around $N_{p}=$ $15\,500\text{\,}$.
@@
All four imputation strategies described in Sec. 5, i.e. mean (MI), block-mean (BMI), hybrid MI-BMI and gappy POD (GPOD) are assessed here. The hybrid mean and block-mean imputation, denoted as $h_{MI-BMI}$, refers to a situation, where a bunch matrix with a standard deviation below unity after the MI is replaced by BMI. Furthermore, the GPOD approach is assessed for three different (fixed) iteration numbers, i.e., GPOD(i), where $i\in[1,5,10]$.
Figure 12 and 13 display two qualitative comparison of the reconstructed jet and the full-order SPH data extracted at three different time instants, with $b=250$, for the u-velocity field and dynamic pressure, respectively.